The diagonal of a rectangle is 17 inches. If the length of the rectangle is 15 inches, what is the width of the rectangle in inches?
Answer by math-vortex(648)

This is a problem where you can use the Pythagorean Equation, a^2 + ^2 = c^2. Sketch a picture of the rectangle and one of its diagonals. Notice that you have a right triangles when you draw the diagonal. The diagonal D is the hypotenuse. The width W and length L of the rectangle form the two legs of the right triangle

The Pythagorean Equation using our variables,
L^2 + W^2 = D^2
.
Substitute in the values we know:
(15)^2 + W^2 = (17)^2
.
Simplify and solve for W. Subtract 225 from both sides.
225 + W^2 = 289
W^2 = 289 – 225
W^2 = 64
.
Take the square root of both sides. If W^2 is 64, then W is 8 because 8 times 8 is 64.
W = 8
.
The width is 8 feet. The problem asks for the width in inches. There are 12 inches in each of those 8 feet, so we multiply 8*12=96. The width is 96 inches.
.
Good luck!
Ms.Figgy